OpenStudy (anonymous):

A 62kg child is sitting on a wagon full of bricks that has a mass of 150kg. In order to move the wagon, the child throws two bricks each of mass 3.0kg in the direction opposite to the direction the wagon is to go. How fast will the wagon move if the bricks are thrown at 2.0ms^-1

OpenStudy (anonymous):

And thats all my questions! thanks(:

OpenStudy (anonymous):

\[P _{1}=P _{2} \] P=mv \[m _{1}v _{1}=m _{2}v _{2}\] 6*2=206*v v=0.056 m/s